A proposed project has fixed costs of $89,000 per year. The operating cash flow at 10,400 units is $127,400. Ignore the effect of taxes.


  

a. What is the degree of operating leverage? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)
b. If units sold rise from 10,400 to 11,100, what will be the new operating cash flow? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
c. What is the new degree of operating leverage? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)
Answer #1

Degree of operating leverage = (fixed costs +operating cash flow)/operating cash flow

=(89,000+10,400)/10,400

= 9.5577

Increase in sales = (11,100-10,400)/10400

= 6.7308%

Hence, increase in operating cash flow = increase in sales *degree of operating leverage

= 6.7308%*9.5577

= 64.33%

Hence, new operating cash flow = 127400+64.33%

=$209,356.42

3. Degree of operating leverage = (209,356.42+89,000)/209,356.42

=$1.4251